OpenStudy (anonymous):

0/2 simplify ? i answer it is undefined and it is still not correct how do i solve?

jimthompson5910 (jim_thompson5910):

remember 0/any number = 0

jimthompson5910 (jim_thompson5910):

you can't have 0/0 though so the general rule is 0/x = 0 where x is any number but 0

OpenStudy (anonymous):

0/2 = 0 2/0 = undefined

OpenStudy (anonymous):

thanks i get it